Find the value of the discriminant b^2-4ac and the number of real solutions for the equatin -x^2+11x-29=0 can not seem to get this one.
Found 2 solutions by stanbon, uma:
Answer by stanbon(75887)

Find the value of the discriminant b^2-4ac and the number of real solutions for the equatin -x^2+11x-29=0 can not seem to get this one.
-----------
b^2-4ac = 11^2-4(-1)(-29)= 5
Therefore, the quadratic has two Real solutions.
------------
x=[-11+-sqrt(5)]/-2
x=(-11/2)+(sqrt5)/2] or x=(-11/2-(sqrt5)/2
Cheers,
Stan H.

Answer by uma(370)

- x^2 + 11x - 29 = 0
Discriminant = b^2 - 4ac
= (11)^2 - 4(-1)(29)
= 22 + 116
= 138
This is positive and not a perfect square.
So the given equation has 2 real and irrational solutions.
Good Luck!!!